# Robustness of topological phases on aperiodic lattices

### v1

@@ 2.4. Position spectral triples and index pairings
A discrete infinite set $\Lambda\subseteq\mathbb{R}^{d}$ is called a Delone set, if there exists positive real numbers $R$ and $r$, such that
#$\#(\operatorname{B}(x,r)\cap\Lambda)\leq 1,\quad\#(\operatorname{B}(x,R)\cap\Lambda)\geq 1$
for all $x\in\mathbb{R}^{d}$, where $\operatorname{B}(x,r)$ denotes the open $r$-ball centered at $x$. We also say the set $\Lambda$ is $(r,R)$-Delone.
@@
Given a Delone set $\Lambda$, equipped with the subspace metric $d(x,y)\mathrel{\vcentcolon=}\left\lvert x-y\right\rvert$. The Voronoi tiling decomposes $\mathbb{R}^{d}$ into tiles lablled by $x\in\Lambda$, and thus decomposes the real Hilbert space $L^{2}(\mathbb{R}^{d})_{\mathbb{R}}$ into a direct sum:
(2.15) $L^{2}(\mathbb{R}^{d})_{\mathbb{R}}\simeq\bigoplus_{x\in\Lambda}L^{2}(T_{x})_{\mathbb{R}}\simeq\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}},$
@@
We may identify all $L^{2}(T_{x})_{\mathbb{R}}$ with a fixed, separable real Hilbert space $\mathcal{K}$. Thus an operator $T\in\mathbb{B}(L^{2}(\mathbb{R}^{d})_{\mathbb{R}})\simeq\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}})$ can be described by its matrix elements $(T_{x,y})_{x,y\in\Lambda}$, where
@@
Definition 2.16.
Let $\Lambda\subseteq\mathbb{R}^{d}$ be a countable discrete subset (e.g. a Delone set). A position spectral triple associated to $\Lambda$ is a real spectral triple of the form
(2.17) $\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
where:
  • $A$ is a real C*-algebra, which carries a *-representation $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}})$ for some separable real Hilbert space $\mathcal{K}$; $\mathcal{A}\subseteq A$ is a dense *-subalgebra;
  • $A\otimes\mathrm{C}\ell_{0,d}$ is represented on $\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ through the tensor product of $\varphi$ and the standard representation of $\mathrm{C}\ell_{0,d}$;
  • ${\scriptstyle\mathsf{X}}_{j}$ is the $j$-th tight-binding “position” operator on $\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}}$, given by
    $({\scriptstyle\mathsf{X}}_{j}f)(x)\mathrel{\vcentcolon=}x_{j}f(x),\qquad x=(x_{1},\dots,x_{d})\in\Lambda\subseteq\mathbb{R}^{d};$
We say a position spectral triple is locally compact, if the matrix elements $\varphi(a)_{x,y}\mathrel{\vcentcolon=}\left\langle x\,\middle|\varphi(a)\middle|\,\mathopen{}y\right\rangle$ are compact for all $a\in\mathcal{A}$ and $x,y\in\Lambda$.
Let $A$ be a real C*-algebra. A position spectral triple of the form (2.17) gives a class in $\operatorname{KK}(A\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Via the Kasparov product, it gives a map
$\operatorname{KK}(\mathrm{C}\ell_{d,0},A)\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
Example 2.18.
Let $\mathbb{R}[\mathbb{Z}^{d}]$ be the real convolution algebra of $\mathbb{Z}^{d}$. That is, the elements of $\mathbb{R}[\mathbb{Z}^{d}]$ are real-valued, finitely supported functions $f\colon\mathbb{Z}^{d}\to\mathbb{R}$. It carries the following convolution product and involution:
The left multiplication of $\mathbb{R}[\mathbb{Z}^{d}]$ on itself extends to a injective *-representation
$\lambda\colon\mathbb{R}[\mathbb{Z}^{d}]\to\mathbb{B}(\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}).$
The closure of $\lambda(\mathbb{R}[\mathbb{Z}^{d}])$ in the real C*-algebra $\mathbb{B}(\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}})$ is the real group C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$.
Let $\mathcal{K}$ be a separable real Hilbert space. Fix a natural number $N$ and represent $\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})$ as well as its dense subalgebra $\mathbb{M}_{N}(\mathbb{R}[\mathbb{Z}^{d}])$ on $\ell^{2}(\mathbb{Z}^{d},\mathcal{K})_{\mathbb{R}}$ using $e_{N}\circ\lambda^{N}$, where $\lambda^{N}$ is the entrywise extension of $\lambda$ to the matrix algebra $\mathbb{M}_{N}\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}=\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathbb{M}_{N}(\mathbb{R})$; and $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\to\mathbb{K}(\mathcal{K})$ is any rank-$N$ corner embedding. Then $\langle f,g\rangle\in\mathbb{M}_{N}(\mathbb{R})$ for all $f,g\in\mathbb{R}[\mathbb{Z}]$. It follows that
$\left(\mathbb{M}_{N}(\mathbb{R}[\mathbb{Z}])\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d},\mathcal{K})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple associated to the Delone set $\Lambda$. It gives a map
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathbb{M}_{N}\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
This spectral triple is used later in Lemma 3.26 from [Ewert-Meyer:Coarse_geometry]*Section 4. It represents the Dirac element of the “real” manifold $\mathbb{T}^{d}$ in the sense of Kasparov [Kasparov:Equivariant_KK-theory]*Definition–Lemma 4.2.
Remark 2.19.
If $\Lambda\subseteq\mathbb{R}^{d}$ is a discrete set but not uniformly discrete, then the Voronoi tiling associated to $\Lambda$ may lack a good physical interpretation. Nevertheless, the above definition still makes sense. This happens, e.g. if we consider the union of two Delone sets $\Lambda_{1}$ and $\Lambda_{2}$, then a position spectral triple associated to $\Lambda\mathrel{\vcentcolon=}\Lambda_{1}\cup\Lambda_{2}$ should be represented on the real Hilbert space $\ell^{2}(\Lambda_{1}\cup\Lambda_{2},\mathcal{K})_{\mathbb{R}}$.
3. Models of aperiodic topological insulators
We describe two observable C*-algebras, modelling a topological insulator that lives on an aperiodic lattice. We assume that the aperiodic lattice is described by a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ (Definition 2.14).
From now on, all Hilbert spaces and C*-algebras are assumed to be real, and we omit their footnotes $\mathbb{R}$ for simplicity.
@@ 3.1. Groupoid C*-algebra of a Delone set
    $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})\subsetneqq\operatorname{UD}_{r}(\mathbb{R}^{d})\subsetneqq\operatorname{QD}(\mathbb{R}^{d})\subsetneqq\mathcal{M}(\mathbb{R}^{d});$
@@
Definition and Lemma 3.2 ([Bourne-Mesland:Aperiodic]*Proposition 2.14).
@@
The groupoid $\mathcal{G}_{\Lambda}$ associated to a Delone set $\Lambda$ is defined as the restriction of $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$ to the transversal $\Omega_{0}$, i.e.
$\mathcal{G}_{\Lambda}\mathrel{\vcentcolon=}\big{(}\Omega_{\Lambda}\rtimes\mathbb{R}^{d}\big{)}^{\Omega_{0}}_{\Omega_{0}}\rightrightarrows\Omega_{0}.$
Then $\mathcal{G}_{\Lambda}$ is an étale groupoid, which is Morita equivalent to $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$.

### v2

@@ 2.3. Position spectral triples
Let $\Lambda\subseteq\mathbb{R}^{d}$ be a countable discrete subset. A position spectral triple associated to $\Lambda$ is a real spectral triple of the form
(2.12) $\xi_{\Lambda}\mathrel{\vcentcolon=}\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad{\scriptstyle\mathsf{X}}\mathrel{\vcentcolon=}\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right),$
where:
  • AA is a real C*-algebra, which carries a *-representation $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda)\otimes\mathcal{K})$ for some separable real Hilbert space $\mathcal{K}$; $\mathcal{A}\subseteq A$ is a dense *-subalgebra;
  • $A\otimes\mathrm{C}\ell_{0,d}$ is represented on $\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ through the tensor product of $\varphi$ and the standard representation of $\mathrm{C}\ell_{0,d}$;
  • ${\scriptstyle\mathsf{X}}_{j}$ is the $j$-th position operator on $\ell^{2}(\Lambda)$, given by
    $({\scriptstyle\mathsf{X}}_{j}\phi)(x)\mathrel{\vcentcolon=}x_{j}\phi(x),\qquad\phi\in\mathrm{C}_{\mathrm{c}}(\Lambda),\;x=(x_{1},\dots,x_{d})\in\Lambda\subseteq\mathbb{R}^{d};$
The condition $\xi_{\Lambda}$ being a real spectral triple says that:
  • $\varphi(a)(1+{\scriptstyle\mathsf{X}}^{2})^{-1}\in\mathbb{K}(\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d})$ for all $a\in\mathcal{A}$;
  • For every $a\in\mathcal{A}$, $\varphi(a)$ maps $\operatorname{dom}{\scriptstyle\mathsf{X}}$ into $\operatorname{dom}{\scriptstyle\mathsf{X}}$, and ${\scriptstyle\mathsf{X}},\varphi(a)]$ extends to a bounded operator on $\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$.
We note the occurence of an auxiliary separable Hilbert space $\mathcal{K}$. The physical meaning of $\mathcal{K}$ is given in Remark˜3.2 for $\Lambda$ a Delone set (cf. Definition˜3.1).
The author thanks the anonymous referee for pointing out the following result:
Lemma 2.13.
Every position spectral triple is locally compact. That is, the matrix elements
$\varphi(a)_{x,y}\mathrel{\vcentcolon=}\langle x\,|\varphi(a)|\,\mathopen{}y\rangle$
are compact for all $a\in\mathcal{A}$ and $x,y\in\Lambda$.
Proof.
By assumption of a real spectral triple, we have that
$\varphi(a)\left(1+\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)^{2}\right)^{-1}\in\mathbb{K}\left(\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}\right)$
for all $a\in\mathcal{A}$. Therefore,
$\left\langle x\,\middle|\varphi(a)\left(1+\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)^{2}\right)^{-1}\middle|\,\mathopen{}y\right\rangle=\left(1+\sum_{j=1}^{d}y_{j}^{2}\right)^{-1}\varphi(a)_{x,y}\otimes\operatorname{id}_{\bigwedge\nolimits^{*}\mathbb{R}^{d}}$
is a compact operator on $\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$. This forces $\varphi(a)_{x,y}\in\mathbb{K}(\mathcal{K})$. ∎
2.3.1. The fundamental class of $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d}$
The last goal of this section is to construct a position spectral triple $\xi_{\mathbb{Z}^{d},N}$ over $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathbb{M}_{N}(\mathbb{R})\otimes\mathrm{C}\ell_{0,d}$, where $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$ is the real group C*-algebra of $\mathbb{Z}^{d}$. Moreover, we shall show that taking the Kasparov product with the class $\xi_{\mathbb{Z}^{d},N}]\in\operatorname{KK}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ gives a surjective homomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
Let $\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d})_{\mathbb{R}}$ be the real convolution algebra of $\mathbb{Z}^{d}$, whose elements are real-valued, finitely supported functions $f\colon\mathbb{Z}^{d}\to\mathbb{R}$, and carries the following convolution product and involution:
The left multiplication of $\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d})_{\mathbb{R}}$ on itself extends to a injective *-representation
$\lambda\colon\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d})_{\mathbb{R}}\to\mathbb{B}(\ell^{2}(\mathbb{Z}^{d},\mathbb{R})).$
The closure of $\lambda(\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d})_{\mathbb{R}})$ in the real C*-algebra $\mathbb{B}(\ell^{2}(\mathbb{Z}^{d},\mathbb{R}))$ is the real group C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$.
Fix a natural number $N$ and extend $\lambda$ entrywise to $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathbb{M}_{N}(\mathbb{R})$. This gives a representation
$\lambda\otimes\operatorname{id}_{N}\colon\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\mathbb{Z}^{d},\mathbb{R})\otimes\mathbb{R}^{N})\simeq\mathbb{B}(\ell^{2}(\mathbb{Z}^{d},\mathbb{R}^{N})).$
To get a position spectral triple over $\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})$, we need another separable real Hilbert space $\mathcal{K}$. Let $\mathbb{R}^{N}\hookrightarrow\mathcal{K}$ be any isometric embedding. Then it induces a corner embedding $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\hookrightarrow\mathbb{K}(\mathcal{K})$ and gives an invertible element in $\operatorname{KK}(\mathbb{M}_{N}(\mathbb{R}),\mathbb{R})$. We represent $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathbb{M}_{N}(\mathbb{R})$ on $\ell^{2}(\mathbb{Z}^{d})\otimes\mathcal{K}$ via $\lambda\otimes e_{N}$. This gives a position spectral triple
(2.14) $\xi_{\mathbb{Z}^{d},N}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d})_{\mathbb{R}})\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d})\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right).$
which represents a class $\xi_{\mathbb{Z}^{d},N}]\in\operatorname{KK}(\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. This is the same class for each $N$ (hence also for $N=1$) because $e_{N}]\in\operatorname{KK}(\mathbb{M}_{N}(\mathbb{R}),\mathbb{R}))$ is invertible.
Theorem 2.15.
Taking the Kasparov product with $\xi_{\mathbb{Z}^{d},N}]$,
$\times_{\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})}[\xi_{\mathbb{Z}^{d},N}]\colon\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}))\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}$
is a surjective group homomorphism.
Proof.
The result is well-known, cf. [Kasparov:K-functor_extension]*Section 5, Theorem 7, whilst we need to appeal to a variant of KK-theory for “real” C*-algebras called KKR-theory (also defined by Kasparov in [Kasparov:K-functor_extension]).
Let $(\mathbb{T}^{d},\tau\colon z\mapsto\overline{z})$ be the “real” $d$-torus. Then Kasparov showed that the “real” spectral triple
(2.16) $\left(\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d})),\;L^{2}(\bigwedge\nolimits^{*}\mathrm{T}^{*}\mathbb{T}^{d}),\;\mathrm{d}+\mathrm{d}^{*}\right)$
represents the KKR-thereotic fundamental class in $\operatorname{KKR}(\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d})),\mathbb{C})$. That is, there exists a class $\beta\in\operatorname{KKR}(\mathbb{C},\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d})))$ (“the Bott class”) such that
$\beta\times_{\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d}))}\alpha=1\in\operatorname{KKR}(\mathbb{C},\mathbb{C}).$
Now we pass from KKR to KK-theory of real C*-algebras. By Lemma˜2.4, the Fourier transform (2.5) maps $\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d}))$ to $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$, where $\mathrm{C}^{*}(\mathbb{Z}^{d})$ carries the real structure $\mathfrak{r}(f)(x)=\overline{f(x)}$; and the Hodge–de Rham operator $\mathrm{d}+\mathrm{d}^{*}$ to the position operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$, which is a “real” operator.
There is a natural isomorphism $\operatorname{KKR}(A,B)\simeq\operatorname{KK}(A^{\mathfrak{r}_{A}},B^{\mathfrak{r}_{B}})$ (cf. [Bourne-Kellendonk-Rennie:Cayley_transform]*Section 2). Under this isomorphism, we find that the Hodge–de Rham spectral triple (2.16), after composing with a rank-one corner embedding $\mathbb{R}\hookrightarrow\mathbb{K}(\mathcal{K})$, is mapped to the position spectral triple $\xi_{\mathbb{Z}^{d},1}$. Hence its Kasparov product with the Bott class in $\operatorname{KK}(\mathbb{R},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d})$ equals $1\in\mathbb{Z}\simeq\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})$. ∎
Remark 2.17.
The KKR-class of the Hodge–de Rham spectral triple (2.16) is also called the Dirac element (cf. [Kasparov:Equivariant_KK-theory]*Definition–Lemma 4.2). It is the KK-theoretic analogue of the fundamental class in cohomology. We refer to [Lord-Rennie-Varilly:Riemannian_manifolds_in_NCG]*Section 4.3 for the notion of fundamental classes in Kasparov theory and noncommutative geometry.
3. Models of aperiodic topological insulators
Convention
From now on, all Hilbert spaces and C*-algebras are assumed to be real. All C*-algebras excepted for the Clifford algebras (and their tensor products with other C*-algebras) are assumed to be ungraded.
In this section, we shall describe the groupoid model (cf. Section˜3.1) and Roe C*-algebra model (cf. Section˜3.2) of topological insulators on aperiodic lattices. A short discussion on the uniform Roe C*-algebras is given in Section˜3.2.3.
We assume that the aperiodic lattice is described by a Delone set $\Lambda\subseteq\mathbb{R}^{d}$, i.e. a uniformly discrete and relatively dense subset of $\mathbb{R}^{d}$, defined as follows.
Definition 3.1.
Let $\Lambda\subseteq\mathbb{R}^{d}$ be a discrete infinite set. Fix $0<r<R$. Then $\Lambda$ is called
  • $r$-uniformly discrete, if $(\operatorname{B}(x,r)\cap\Lambda)\leq 1$ for all $x\in\mathbb{R}^{d}$;
  • $R$-relatively dense, if $(\operatorname{B}(x,R)\cap\Lambda)\geq 1$ for all $x\in\mathbb{R}^{d}$;
  • $(r,R)$-Delone, if $\Lambda$ is both $r$-uniformly discrete and $R$-relatively dense.
@@ 3. Models of aperiodic topological insulators
Remark 3.2.
Given a Delone set $\Lambda$, equipped with the subspace metric $d(x,y)\mathrel{\vcentcolon=}\left\lvert x-y\right\rvert$. The Voronoi tiling (cf. Figure˜3.1) decomposes $\mathbb{R}^{d}$ into tiles lablled by $x\in\Lambda$, and thus decomposes the Hilbert space $L^{2}(\mathbb{R}^{d})$ into a direct sum:
(3.3) $L^{2}(\mathbb{R}^{d})\simeq\bigoplus_{x\in\Lambda}L^{2}(T_{x}),$
@@
We may further identify all $L^{2}(T_{x})$’s with a fixed, separable Hilbert space $\mathcal{K}$. Thus an operator $T\in\mathbb{B}(L^{2}(\mathbb{R}^{d}))\simeq\mathbb{B}(\ell^{2}(\Lambda)\otimes\mathcal{K})$ can be described by its matrix elements $(T_{x,y})_{x,y\in\Lambda}$, where
@@
Figure 3.1. Voronoi tiling associated to a Delone set
@@ 3.1. Groupoid C*-algebra of a Delone set
    $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})\subsetneqq\operatorname{UD}_{r}(\mathbb{R}^{d})\subsetneqq\operatorname{QD}(\mathbb{R}^{d})\subsetneqq\mathcal{M}(\mathbb{R}^{d})\textup{;}$
@@
Definition and Lemma 3.5 ([Bourne-Mesland:Topological_phases]*Proposition 2.14).
@@
Define the groupoid $\mathcal{G}_{\Lambda}$ as the restriction of $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$ to the transversal $\Omega_{0}$. Then $\mathcal{G}_{\Lambda}$ is an étale groupoid, which is Morita equivalent to $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$.